OpenStudy (anonymous):

use the method of substitution to solve the system of linear equation. 5a=2-b 5a=7-b A. there is one solution, the solution is___ B. there are infinitely many solutions C. there are no solutions.

OpenStudy (anonymous):

You can just look at this and see that 5a cannot be equal to 2 - b and also 7 - b at the same time. Right?

OpenStudy (anonymous):

Can you think of a b that works? Or 7-b = 2-b b = ?

OpenStudy (anonymous):

I just learned how to do this

OpenStudy (anonymous):

7-2=2-7?

OpenStudy (anonymous):

Answer C.

OpenStudy (anonymous):

yes